# Spherical maximal functions and Hardy spaces for Fourier integral operators

### v2

@@ 4.2. Sharpness for hypersurfaces
(4.16) $\mathcal{C}_{1}:=\{(x,t,\partial_{x}\Phi(z,\eta),\partial_{t}\Phi(z,\eta),\partial_{\eta}\Phi(x,t,\eta),\eta)\mid(x,t,\eta)\in V\}.$
On the other hand, since $R$ is smoothing, the canonical relation of $TS$ coincides with that of $\chi_{1}\mathcal{A}S_{2}$, which in turn is contained in a small neighborhood of $c_{0}$ in $\mathcal{C}$. Moreover, since $S$ is a change of coordinates on $\mathbb{R}^{n}$, this means that the canonical relation of $T$ is contained in
@@
Next, we consider Theorem 3.9, which we show to be essentially sharp for all $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty]$ whenever the function $\rho$ is not identically under natural nontriviality assumptions. We use notation as in Theorem 3.9.
@@
We also obtain sharpness for all $\alpha\in\mathbb{C}$ and $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty]$ in Proposition 3.11, as a consequence of the following result. Recall the definition of $\mathcal{M}_{t}^{\alpha}$ from (3.22), for $\alpha\in\mathbb{C}$ and $t>0$.
@@
for all $\xi\in\mathbb{R}^{n}$, where $a_{\alpha}^{0}\in C^{\infty}_{c}(\mathbb{R}^{n})$ and $a_{\alpha}^{+},a_{\alpha}^{-}\in S^{-\frac{n-1}{2}-\operatorname{Re}(\alpha)}(\mathbb{R}^{n})$. In fact, it follows from asymptotics of Bessel functions (see [38, Section 7.21]) that $a^{+}_{\alpha}$ and $a^{-}_{\alpha}$ are elliptic. For $t>0$ and $\xi\in\mathbb{R}$, set
$T_{t}^{+}(\xi):=e^{it|\xi|}a_{\alpha}^{+}(t\xi)\ \text{ and }\ T_{t}^{-}(\xi):=e^{-it|\xi|}a_{\alpha}^{-}(t\xi)$
@@
Now, let $t_{0}>0$ be as in (4.18). Fix $x_{0}\in\mathbb{R}^{n}$, and let $S_{1}$ and $S_{2}$ be pseudodifferential operators of order zero, with principal symbols that are nonzero at $(x_{0},e_{1})$ and $(x_{0}+t_{0}e_{1},e_{1})$, respectively, and with microsupport contained in small neighborhoods of these points. Here $e_{1}\in\mathbb{R}^{n}$ is the first basis vector. Then, by construction, the principal symbol of $S_{1}T_{t_{0}}^{+}(D)S_{2}$ is nonzero at $(x_{0},e_{1},x_{0}+t_{0}e_{1},e_{1})$, and the canonical relation of $S_{1}T_{t_{0}}^{+}(D)S_{2}$ is contained in a small conic neighborhood in $\mathcal{C}^{+}_{t}$ of this point. Moreover, the canonical relation of $S_{1}T_{t_{0}}^{-}(D)S_{2}$ is empty, so $S_{1}T_{t_{0}}^{-}(D)S_{2}$ is a smoothing operator.
By (4.18) and because $L^{p}(\mathbb{R}^{n})$ and $\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ are invariant under pseudodifferential operators of order zero, one has $\|S_{1}\mathcal{M}_{t_{0}}^{\alpha}S_{2}f\|_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$ for all $f\in\mathcal{S}(\mathbb{R}^{n})$. In turn, since $a_{\alpha}^{0}(t_{0}D)$ and $S_{1}T_{t_{0}}^{-}(D)S_{2}$ are smoothing operators, this implies that
$\|S_{1}T_{t_{0}}^{+}(D)S_{2}f\|_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$
for $f\in\mathcal{S}(\mathbb{R}^{n})$. Moreover, one can again write $S_{1}T_{t_{0}}^{+}(D)S_{2}$ as the sum of an operator as in Theorem 4.2 precomposed with a change of coordinates, and a harmless smoothing operator. So Theorem 4.2 shows that $s\geq s(p)-\frac{n-1}{2}-\operatorname{Re}(\alpha)$ if $p\geq 2$.
Next, let $I\subseteq(0,\infty)$ be a non-empty open interval such that (4.19) holds, and fix $t_{0}\in I$ and $x_{0}\in\mathbb{R}^{n}$. Let $S_{2}$ be as before, and let $\chi_{1}\in C^{\infty}_{c}(\mathbb{R}^{n+1})$ be a smooth cutoff which localizes to a small neighborhood of $(x_{0},t_{0})$. Then (4.19) implies that
$\big{\|}\sup_{t\in\mathbb{R}}|\chi_{1}(\cdot,t)\mathcal{M}_{t}^{\alpha}S_{2}f(\cdot)|\big{\|}_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$
for all $f\in\mathcal{S}(\mathbb{R}^{n})$. Set $T^{+}f(x,t):=T_{t}^{+}(D)f(x)$ and $T^{-}f(x,t):=T^{-}_{t}(D)f(x)$ for $x\in\mathbb{R}^{n}$ and $t>0$. Then $S_{1}T^{-}S_{2}$ is a smoothing operator, as is the term corresponding to $a_{\alpha}^{0}$. It then follows that
$\big{\|}\sup_{t\in\mathbb{R}}|\chi_{1}(\cdot,t)T^{+}S_{2}f(\cdot,t)|\big{\|}_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$
for all $f\in\mathcal{S}(\mathbb{R}^{n})$. Since the the canonical relation of $\chi_{1}T^{+}S_{2}$ is contained in
@@ 4.3. Sharpness on manifolds
We first observe that the maximal function estimate in Proposition 3.14 (see also Remark 3.3) is essentially sharp for all $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty]$.
@@
for all $f\in\mathcal{D}(M,\Omega_{1/2})$, then $s\geq\frac{n-1}{2}(\frac{1}{2}-\frac{1}{p})$. Moreover, if there exists a non-empty open interval $I\subseteq\mathbb{R}$ and a $C\geq 0$ such that
@@
For the first statement one can reduce to Theorem 4.2 by working in local coordinates. Alternatively, one can reason as follows. By replacing $f$ by $e^{-it_{0}\sqrt{-\Delta_{g}}}f$ and using the invariance of $\mathcal{H}^{s,p}_{FIO}(M)$ under $e^{-it_{0}\sqrt{-\Delta_{g}}}$, we may suppose that $t_{0}=0$. But then the first statement follows from the fact that the Sobolev embeddings in (2.11) are sharp (which in turn is a consequence of, e.g., Theorem 4.2).
@@
Next, one can argue as in the proof of Theorem 3.6 to express $T$ microlocally in terms of an operator as in (4.1), with $m=0$, after which one can combine (4.21) with Theorems 4.2 and 4.5. This is possible because $T$ is a Fourier integral operator of order $-1/4$, with a principal symbol which does not vanish identically, and with a canonical relation of the form
(4.22) $\mathcal{C}:=\{(x,t,\xi,-p(x,\xi),y,\eta)\mid(x,\xi)=\chi_{t}(y,\eta)\}.$
Here $\chi_{t}$ is the bicharacteristic flow at time $t$ associated with $\sqrt{g}$ in local coordinates, and $p(x,\xi)>0$. These statements about $T$ in turn follow by expressing the canonical relation of $(e^{it\sqrt{-\Delta_{g}}})_{t\in\mathbb{R}}$ (see e.g. [17, Theorem 29.11]) in local coordinates. In particular, as in the proof of Theorem 4.8, the fact that $p$ is nonzero implies that the resulting operator satisfies the conditions of Theorem 4.5. ∎
Finally, the maximal function estimate in Proposition 3.15 is also essentially sharp for all $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty]$. We use notation as in (3.28).
@@
Let $p\in[1,\infty]$ and $s\in\mathbb{R}$. If if there exist $t_{0}\in(0,\mathrm{inj}(M))$ and $C_{t_{0}}\geq 0$ such that
@@
for all $f\in\mathcal{D}(M,\Omega_{1/2})$, then $s\geq\frac{n-1}{2}(\frac{1}{2}-\frac{1}{p})-\frac{n-1}{2}$. Moreover, if if there exists a non-empty open interval $I\subseteq(0,\mathrm{inj}(M))$ and a $C\geq 0$ such that
@@
Firstly, by working in local coordinates as in the proof of Proposition 4.12, one can reduce to an application of Theorems 4.2 and 4.5. This is possible for the first statement because $\mathcal{A}_{t_{0},g}$ is a Fourier integral operator of order $-(n-1)/2$, associated with a local canonical graph and with a non-vanishing principal symbol. For the second statement one can also use that the canonical relation of the operator $T$, given by $Tf(x,t):=\chi(t)\mathcal{A}_{t,g}f(x)$ for $f\in\mathcal{D}(M,\Omega_{1/2})$, $x\in M$, $t\in\mathbb{R}$ and $0\neq\chi\in C^{\infty}_{c}(\mathbb{R})$ such that $\mathrm{supp}(\chi)\subseteq I$, is contained in the union of the canonical relation in (4.22) and the corresponding one with $-p(x,\xi)$ replaced by $p(x,\xi)$.

### v3

@@ 4.2. Sharpness for hypersurfaces
(4.16) $\mathcal{C}_{1}:=\{(x,t,\partial_{x}\Phi(x,t,\eta),\partial_{t}\Phi(x,t,\eta),\partial_{\eta}\Phi(x,t,\eta),\eta)\mid(x,t,\eta)\in V\}.$
On the other hand, since $R$ is smoothing, the canonical relation of $TS^{\prime}$ coincides with that of $\chi_{1}\mathcal{A}S$, which in turn is contained in a small neighborhood of $c_{0}$ in $\mathcal{C}$. Moreover, since $S^{\prime}$ is a change of coordinates on $\mathbb{R}^{n}$, this means that the canonical relation of $T$ is contained in
@@
Next, we consider Theorem 3.10, which we show to be essentially sharp for all $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty)$ under natural nontriviality assumptions (see also Remark 3.4 for $p=\infty$). We use notation as in Theorem 3.10.
@@
We also obtain sharpness for all $\alpha\in\mathbb{C}$ and $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty)$ in Proposition 3.12 (see Remark 3.4 for $p=\infty$), as a consequence of the following result. Recall the definition of $\mathcal{M}_{t}^{\alpha}$ from (3.25), for $\alpha\in\mathbb{C}$ and $t>0$.
@@
for all $\xi\in\mathbb{R}^{n}$, where $a_{\alpha}^{0}\in C^{\infty}_{c}(\mathbb{R}^{n})$ and $a_{\alpha}^{+},a_{\alpha}^{-}\in S^{-\frac{n-1}{2}-\operatorname{Re}(\alpha)}(\mathbb{R}^{n})$. In fact, it follows from asymptotics of Bessel functions (see [43, Section 7.21]) that $a^{+}_{\alpha}$ and $a^{-}_{\alpha}$ are elliptic. For $t>0$ and $\xi\in\mathbb{R}$, set
[…]
@@
Now, let $t_{0}>0$ be as in (4.18). Fix $x_{0}\in\mathbb{R}^{n}$, and let $S$ be a pseudodifferential operator of order zero, with a principal symbol which is nonzero at $(x_{0}+t_{0}e_{1},e_{1})$, and with microsupport contained in a small neighborhoods of this point. Here $e_{1}\in\mathbb{R}^{n}$ is the first basis vector. Then, by construction and because $\mathcal{C}_{t}^{+}$ can be parametrized using its last two coordinates, the principal symbol of $T_{t_{0}}^{+}(D)S$ is nonzero at $(x_{0},e_{1},x_{0}+t_{0}e_{1},e_{1})$, and the canonical relation of $T_{t_{0}}^{+}(D)S$ is contained in a small conic neighborhood in $\mathcal{C}^{+}_{t}$ of this point. Moreover, the canonical relation of $T_{t_{0}}^{-}(D)S$ is empty, so $T_{t_{0}}^{-}(D)S$ is a smoothing operator.
By (4.18) and because $\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ is invariant under pseudodifferential operators of order zero, one has $\|\mathcal{M}_{t_{0}}^{\alpha}Sf\|_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$ for all $f\in\mathcal{S}(\mathbb{R}^{n})$. In turn, since $a_{\alpha}^{0}(t_{0}D)$ and $T_{t_{0}}^{-}(D)S$ are smoothing operators, this implies that
$\|T_{t_{0}}^{+}(D)Sf\|_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$
for $f\in\mathcal{S}(\mathbb{R}^{n})$. Moreover, one can again write $T_{t_{0}}^{+}(D)S$ as the sum of an operator as in Theorem 4.2 precomposed with a change of coordinates, and a harmless smoothing operator. So Theorem 4.2 shows that $s\geq s(p)-\frac{n-1}{2}-\operatorname{Re}(\alpha)$ if $p\geq 2$.
Next, let $I\subseteq(0,\infty)$ be a non-empty open interval such that (4.19) holds, and fix $t_{0}\in I$ and $x_{0}\in\mathbb{R}^{n}$. Let $S$ be as before, and let $\chi\in C^{\infty}_{c}(\mathbb{R}^{n+1})$ be a smooth cutoff which localizes to a small neighborhood of $(x_{0},t_{0})$. Then (4.19) implies that
$\big\|\sup_{t\in\mathbb{R}}|\chi(\cdot,t)\mathcal{M}_{t}^{\alpha}Sf(\cdot)|\big\|_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$
for all $f\in\mathcal{S}(\mathbb{R}^{n})$. Set $T^{+}f(x,t):=T_{t}^{+}(D)f(x)$ and $T^{-}f(x,t):=T^{-}_{t}(D)f(x)$ for $x\in\mathbb{R}^{n}$ and $t>0$. Then $T^{-}S$ is a smoothing operator, as is the term corresponding to $a_{\alpha}^{0}$. It thus follows that
$\big\|\sup_{t\in\mathbb{R}}|{\chi(\cdot,t)T^{+}Sf(\cdot,t)}|\big\|_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})}$
for all $f\in\mathcal{S}(\mathbb{R}^{n})$. Since the the canonical relation of $\chi T^{+}S$ is contained in
@@ 4.3. Sharpness on manifolds
We first observe that the maximal function estimate in Proposition 3.15 (see also Remark 3.4) is essentially sharp for all $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty)$.
@@
for all $f\in\mathcal{D}(M,\Omega_{1/2})$, then $s\geq\frac{n-1}{2}(\frac{1}{2}-\frac{1}{p})$. Moreover, if there exist a non-empty open interval $I\subseteq\mathbb{R}$ and a $C\geq 0$ such that
@@
For the first statement one can reduce to Theorem 4.2 by working in local coordinates. Alternatively, one can reason as follows. By replacing $f$ by $e^{-it_{0}\sqrt{-\Delta_{g}}}f$ and using the invariance of $\mathcal{H}^{s,p}_{FIO}(M)$ under $e^{-it_{0}\sqrt{-\Delta_{g}}}$, we may suppose that $t_{0}=0$. But then the first statement follows from the fact that the Sobolev embeddings in (2.11) are sharp (which in turn is a consequence of e.g. Theorem 4.2).
@@
Next, one can argue as in the proof of Theorem 3.7 to express $T$ microlocally in terms of an operator as in (4.1), with $m=0$, after which one can combine (4.21) with Theorems 4.2 and 4.5. This is possible because $T$ is a Fourier integral operator of order 1/4-1/4, with a principal symbol which does not vanish identically, and with a canonical relation of the form
(4.22) $\mathcal{C}:=\{(x,t,\xi,p(x,\xi),y,\eta)\mid(x,\xi)=\chi_{t}(y,\eta)\}.$
Here $\chi_{t}$ is the bicharacteristic flow at time $t$ associated with $\sqrt{g}$ in local coordinates, and $p(x,\xi)>0$ is the principal symbol of $\sqrt{-\Delta_{g}}$ in local coordinates. These statements about $T$ in turn follow by expressing the canonical relation of $(e^{it\sqrt{-\Delta_{g}}})_{t\in\mathbb{R}}$ (see e.g. [19, Theorem 29.1.1]) in local coordinates. In particular, as in the proof of Theorem 4.8, the fact that the symbol $p$ is nonzero implies that the resulting operator satisfies the conditions of Theorem 4.5. ∎
Finally, the maximal function estimate in Proposition 3.16 is also essentially sharp for all $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty)$ (see Remark 3.4 for $p=\infty$). We use notation as in (3.32).
@@
Let $p\in[1,\infty]$ and $s\in\mathbb{R}$. If there exist $t_{0}\in(0,\mathrm{inj}(M))$ and $C_{t_{0}}\geq 0$ such that
@@
for all $f\in\mathcal{D}(M,\Omega_{1/2})$, then $s\geq\frac{n-1}{2}(\frac{1}{2}-\frac{1}{p})-\frac{n-1}{2}$. Moreover, if there exist a non-empty open interval $I\subseteq(0,\mathrm{inj}(M))$ and a $C\geq 0$ such that
@@
Firstly, by working in local coordinates as in the proof of Proposition 4.12, one can reduce to an application of Theorems 4.2 and 4.5. This is possible for the first statement because $\mathcal{A}_{t_{0},g}$ is a Fourier integral operator of order (n1)/2-(n-1)/2, associated with a local canonical graph and with a non-vanishing principal symbol. For the second statement one can also use that the canonical relation of the operator $T$, given by $Tf(x,t):=\chi(t)\mathcal{A}_{t,g}f(x)$ for $f\in\mathcal{D}(M,\Omega_{1/2})$, $x\in M$, $t\in\mathbb{R}$ and $0\neq\chi\in C^{\infty}_{c}(\mathbb{R})$ such that $\mathrm{supp}(\chi)\subseteq I$, is contained in the union of the canonical relation in (4.22) and the corresponding one with $p(x,\xi)$ replaced by $p(x,\xi)$.
@@
Acknowledgments
The authors would like to thank the referee for carefully reading the manuscript and for several very helpful suggestions.